No number

Proposal for a L^{A}TEX document class for SAE Technical Papers

Axel Franke

Lund Institute of Technology, Sweden

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999 Society of Automotive Engineers, Inc.

ABSTRACT

This L^{A}TEX class provides correct formatting according to the requirements for the publication of an SAE Technical Paper. It is supposed to be useful for authors publishing simulation models and similar containing lots of formulas. The style contains many features automating the formatting and thus, making it easier for the author to concentrate on contents.

INTRODUCTION

This is a Level 1 heading, created with the command

\section{Introduction}

As can be seen, the correct format (bold face and capitalized) is done by the style – the author does not need to care about that.

TITLE AND AUTHOR INFORMATION

Title, author information, paper number and copyright are created in the preface using the commands

\PaperTitle{Proposal for a \LaTeX\ document
class for SAE Technical Papers}
\AddAuthor{Axel Franke}{Lund Institute of
Technology, Sweden}
\PaperNumber{No number}
\SAECopyright{1999}

Authors from the same affiliation should be added within one \AddAuthor command. The general syntax is

\AddAuthor{Author1, Author2, ...}{Affiliation}

The command may be used several times.

A SUBHEADING This is a subheading which has been created as usual with the command

\subsection{A subsection}

Also the following sub-subheading is formatted automatically by the style.

\underline{General options of the style} The sae L^{A}TEX style supports optional arguments which determine the document font (Helvetica or Times) and the paper size (letterpaper or A4). Default is Helvetica and A4. The letter format of this publication has been obtained by the opening command

                                                       \documentclass[letterpaper]{sae}

If Times is desired as document font, this can be achieved by using

                                                       \documentclass[letterpaper,times]{sae}

or

                                                       \documentclass[times]{sae}

The latter uses A4 as paper size. This concept has the advantage that document font and the paper size can be changed easily at any time of the publication process.